OpenStudy (anonymous):

what is the correct classification of the polynomial 5x^3y^3+2x^2y^3+x^2y^2+3xy^2

OpenStudy (kc_kennylau):

Order 6, if you're asking about order :) (just guessing)

OpenStudy (anonymous):

\[5x ^{3}y ^{3}+2x ^{2}y ^{3}+x ^{2}y ^{2}+3xy ^{2}\]

OpenStudy (anonymous):

okay how do u get the order?

OpenStudy (kc_kennylau):

Degree of first term: \(3+3=6\) Degree of second term: \(2+3=5\) Degree of third term: \(2+2=4\) Degree of last term: \(1+2=3\) Highest degree: 6 \(\therefore\) order=6 (To get the degree of a term, add the indices together :p)

OpenStudy (anonymous):

the options given are a. A 4-term polynomial in x and y of degree 3 b. A 4-term polynomial in x and y of degree 6 c. A 4-term polynomial in x and y of degree 5 d. A 4-term polynomial in x and y of degree 2

OpenStudy (anonymous):

okay so the answer is b right? :D

OpenStudy (kc_kennylau):

yes :)

OpenStudy (kc_kennylau):

If you wanna practice, find out the degree of \(xyz+3x^2y\) :)

OpenStudy (anonymous):

the degree is 4, right? if z is also included

OpenStudy (anonymous):

sorry i mean 3 not 4

OpenStudy (kc_kennylau):

yes :D

OpenStudy (anonymous):

alright thank you so much :)

OpenStudy (kc_kennylau):

no problem :)